• Main INDEX
  • Monthly INDEX
  • NEXT
    Make New Entry, Make Followup Entry

    User name norum

    Log entry time 00:07:33 on March 1,2011

    Entry number 349283

    keyword=Shift summary

    16:00 Beam permit
    E08-10 Kinematics # 1 (LHRS-0.552 GeV/c, 25.5 deg;
    RHRS-0.805, 12.5)
    16:01 Run 2051 started - Cosmics
    16:18 Run 2052 started - Cosmics
    16:28 Beam check run started.
    17:05 Running at 20 uA still gave >20% DT. BigBite and Hand
    removed from DAQ. Configuration changed to "d Threshold - 2
    Arm"
    17:12 Production running at 15 uA and DT=12-13% on 4 cm LH2
    Run 2059 started - Production
    17:47 Run 2060 started - Production
    18:34 Run 2061
    18:55 All FPP chamber HV's ramped to zero due to lack of CO2.
    New CO2 tank being attached; C. Hanretty will reset VH in
    morning.
    19:00 Small blip in LHRS HV's (Chassis 7) to be expected due to
    wrong icon being pinged then immediately reset.
    19:19 Run 2062
    20:04 Run 2063
    20:36 Run 2064
    21:20 Run 2065
    21:45 Escorted access to remove detector S2 from the trigger in RHRS
    22:53 Beam requested
    23:15 Run 2066 - Production
    23:49 Run 2067